In 2017, the rate of job separations was 0.015 and the rate of job findings was 0.185. In 2018 the rates are 0.07 and 0.28, respectively. According to the natural rate of unemployment, what is the percent change in the unemployment rate between 2017 and 2018?

Group of answer choices

27.5 percent

166.7 percent

30 percent

22.7 percent

15 percent

33.3 percent

44.3 percent

25 percent
Answer #1

Natural rate of unemployment = (rate of job separation) / (rate of job finding + rate of job separation)

1) Natural rate of unemployment 2017 = (0.015)/(0.185 + 0.015) = 7.5%

2) Natural rate of unemployment 2018 = (0.07)/(0.07 + 0.28) = 20.0%

Change in the rate of unemployment = (20 - 7.5)/7.5 = 166.67

Select 166.67 percent